for 4-methylbenzaldehyde can someone explain all the splitting patterns present in the mass spectrum, aka why they are there(such as m+ peak, doublets,etc) thank you

Int(120) 20- 25 50 75 100 125
Answer #1

In mass spectrum we are not getting splitting pattern we getting ions ,in nmr splitting is take place because of adjacent hydrogen atoms.

CHO CH 7 rol Lo ін- CH3 m/2 = 120 m/2=119 СН3 m/2 = 9 119 are obtaines m2 = 120 and m/2 = by loss of H-atom